# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 1. $\blacksquare_{1}$
    for $D\in\{C_{1}^{*},L^{*},C_{2}^{*},K^{*},C_{3}^{*}\})\ \pi^{*}[D]=h^{*}[D]$, and $\pi^{*}$, $h^{*}$ map endpoints of $D$ to endpoints of $\pi^{*}[D]$ (which is $C_{1}^{*}$ if $D=C_{1}^{*}$, etc.),
@@ 3. $\blacksquare_{3}$
    in particular, if $D\neq E\in\{C_{1}^{*},L^{*},C_{2}^{*},K^{*},C_{3}^{*}\}$, then $\pi^{*}[D]\cap\pi^{*}[E]=\emptyset$,
Next, we define an extension $J^{**}$ of $J^{*}$ as follows. The pair $J^{**}$ and $\pi^{**}\in{\rm Epi}(J^{**},J^{*})$ is in fact uniquely determined (up to isomorphism) by the requirements that
  • $J^{**}$ is the disjoint union of $D_{i}$, $i=1,2,\ldots,8$ with $D_{i}RD_{i+1}$,
  • for each $i$ the map $\pi^{**}{\restriction}D_{i}$ is a bijection,
  • $\pi^{**}[D_{1}]=\pi^{**}[D_{8}]=J^{*}$,
  • $\pi^{**}[D_{2}]=\pi^{**}[D_{7}]=C_{3}^{*}$,
  • $\pi^{**}[D_{i}]=K^{*}$ for $i=3,4,5,6$.
@@
(5.23) $(\pi^{**}{\restriction}D_{8}{\textrm{ is an isomorphism with }}J^{*})\ \&\ (d\in D_{8}\wedge dRD_{7}\ \to\pi^{**}(d)\in C_{3}^{*}),$
(5.24) $(\pi^{**}{\restriction}D_{1}{\textrm{ is an isomorphism with }}J^{*})\ \&\ (d\in D_{1}\wedge dRD_{2}\ \to\pi^{**}(d)\in C_{3}^{*}),$
Next, we are going to use $M$ and $f$ from $(7)$ to construct $J^{\prime\prime}$. W.l.o.g. we can assume that $f$ factors through $\pi^{*}\circ\pi^{**}$, or simply $f$ maps onto $J^{**}$, so we can pick $J^{\prime\prime}$, $f^{*}\in{\rm Epi}(J^{\prime\prime},M)$, $\pi^{\prime\prime}\in{\rm Epi}(J^{\prime\prime},J^{**})$ with $\pi^{\prime\prime}=f\circ f^{*}$. We can assume that $(\pi^{\prime\prime})^{-1}(a)$ consists of intervals each of length at least $3$ whenever $a\in J^{**}$.
Finally, we claim that letting $\pi^{\prime}=\pi^{*}\circ\pi^{**}\circ\pi^{\prime\prime}$, $h=h^{*}\circ\pi^{**}\circ\pi^{\prime\prime}$, and $K^{\prime\prime}$ be any interval in $J^{\prime\prime}$ with
  • $\pi^{\prime\prime}[K^{\prime\prime}]=D_{4}\cup D_{5}$, and $K^{\prime\prime}$ is maximal such interval (i.e. $a\notin K^{\prime\prime}$, $aRK^{\prime\prime}$ implies $\pi^{\prime\prime}(a)\in D_{3}\cup D_{6}$),
@@
will work (i.e. $\boxplus_{6}$ is satisfied). We also remark that passing to a further extension of $J^{\prime\prime}$, if necessary and replacing $J^{\prime\prime}$, $\pi^{\prime\prime}$, without loss of generality such $K^{\prime\prime}$ exists.
@@
It is easy to see that $\pi^{\prime\prime}(c)\in D_{3}\cup D_{6}$. We let $k^{*}_{-}$ denote the unique node in $K^{*}$ with $k^{*}_{-}RC_{2}^{*}$. By the way $\pi^{**}:J^{**}\to J^{*}$ is defined one can check that
$\pi^{**}\circ\pi^{\prime\prime}(c)=k^{*}_{-}.$
@@
(5.25) ${\rm tp}^{c,C}(\pi^{**}\circ\pi^{\prime\prime})\supseteq\{\{k^{*}_{-}\},K^{*},K^{*}\cup C_{3}^{*},K^{*}\cup C_{3}^{*}\cup C_{2}^{*},K^{*}\cup C_{3}^{*}\cup C_{2}^{*}\cup L^{*}\}.$
Clearly ${\rm tp}^{c,C}(\pi^{\prime\prime})$ contains a set $H$, such that $H$ is either the entire $D_{3}$ and a subset of $D_{4}\cup D_{5}\cup D_{6}$, or $H$ contains the entire $D_{6}$, and a subset of $D_{3}\cup D_{4}\cup D_{5}$. Either case $\pi^{**}[H]=K^{*}$. Similarly, ${\rm tp}^{c,C}(\pi^{\prime\prime})$ contains a set $H^{\prime}$ that is either the union of $D_{2}\cup D_{3}$ and a subset of $D_{4}\cup D_{5}\cup D_{6}\cup D_{7}$, or $H$ is the union of the full $D_{6}\cup D_{7}$ and a subset $D_{2}\cup D_{3}\cup D_{4}\cup D_{5}$. In either case, $\pi^{**}[H]=K^{*}\cup C_{3}^{*}$. This implies that
${\rm tp}^{c,C}(\pi^{**}\circ\pi^{\prime\prime})\supseteq\{\{k^{*}_{-}\},K^{*},K^{*}\cup C_{3}^{*}\}.$
Now recalling (5.23) and (5.24) it is easy to check that (5.25) holds, indeed. Finally, $k^{*}_{-}$ is an endpoint of $K^{*}$, so by $\blacksquare_{1}$ and $\blacksquare_{3}$ the node $\pi^{*}(k^{*}_{-})=\pi^{\prime}(c)=h^{*}(k^{*}_{-})=h^{\prime}(c)$ is an endpoint of $K^{\prime}$, so applying $\pi^{*}$ and $h^{*}$ to (5.25),
${\rm tp}^{c,C}(\pi^{*}\circ\pi^{**}\circ\pi^{\prime\prime})\supseteq\{\{\pi^{\prime}(k^{\prime\prime}_{-})\},K^{\prime},K^{\prime}\cup C_{3}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime}\cup L^{\prime}\},$
@@
${\rm tp}^{c,C}(h^{*}\circ\pi^{**}\circ\pi^{\prime\prime})\supseteq\{\{\pi^{\prime}(k^{\prime\prime}_{-})=h^{*}(k^{*}_{-})\},K^{\prime},K^{\prime}\cup C_{3}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime}\cup L^{\prime}\},$
which by our assumption $(5)$ on the structure of $J^{\prime}$ uniquely determine ${\rm tp}^{c,C}(\pi^{*}\circ\pi^{**}\circ\pi^{\prime\prime})$ and ${\rm tp}^{c,C}(h^{*}\circ\pi^{**}\circ\pi^{\prime\prime})$, in particular these two coincide. ∎
Appendix A Projective Fraïssé limits
We recall here the framework of projective Fraïssé theory. The theory was introduced in [7]. The description below is a generalized version of [7].
@@ 2. $({\rm G}2)$
be the quotient map, which we call the projection.
Fact A.2.
If $x\in{\mathbb{K}}$, then sets of the form
@@
Suppose that $U\subseteq K$ is open and ${\rm pr}(x)\in U$. Since there are only countably many epimorphisms in $\mathcal{K}$}, by the amalgamation property of $\mathcal{K}$} (and by $({\rm P}1)$, $({\rm P}2)$), we can enumerate a cofinal system $(\varphi_{i})_{i}$ of epimorphisms, in the sense that each epimorphism factors through all, but finitely many $\varphi_{i}$. For each $\varphi_{i}$ pick $y_{i}\in{\mathbb{K}}$ with ${\rm pr}(y_{i})\notin U$, $\varphi(y_{i})R\varphi(x)$. By the previous remark about $(\varphi_{i})_{i}$ (i.e. for each $i\leq j$ there exists $\pi$ such that $\varphi_{i}=\pi\circ\varphi_{j}$) and by a standard compactness argument we can assume $\varphi_{i}(y_{j})=\varphi_{i}(y_{i}))$ for $i\leq j$. By compactness of $\mathbb{K}$}, we can assume that $(y_{i})_{i}$ converges to some $y\in{\mathbb{K}}$ (in fact the convergence of $(\varphi_{i}(y_{j}))_{j}$ already implies this), so by continuity,
$\textrm{for each }i,\ \varphi_{i}(y)=\varphi_{i}(\lim_{j}(y_{j}))=\lim_{j}(\varphi_{i}(y_{j}))=\varphi_{i}(y_{i}),$

### v2

@@ 1. $\blacksquare_{1}$
    for $D\in\{C_{1}^{+},L^{+},C_{2}^{+},K^{+},C_{3}^{+}\})\ \pi^{+}[D]=h^{+}[D]$, and $\pi^{+}$, $h^{+}$ map endpoints of $D$ to endpoints of $\pi^{+}[D]$ (which is $C_{1}^{+}$ if $D=C_{1}^{+}$, etc.),
@@ 3. $\blacksquare_{3}$
    in particular, if $D\neq E\in\{C_{1}^{+},L^{+},C_{2}^{+},K^{+},C_{3}^{+}\}$, then $\pi^{+}[D]\cap\pi^{+}[E]=\emptyset$,
Next, we define an extension $J^{++}$ of $J^{+}$ as follows. The pair $J^{++}$ and $\pi^{++}\in{\rm Epi}(J^{++},J^{+})$ will be an unfolding of $J^{+}$, which is in fact uniquely determined (up to isomorphism) by the requirements that
  • $J^{++}$ is the disjoint union of $D_{i}$, $i=1,2,\ldots,8$ with $D_{i}RD_{i+1}$,
  • for each $i$ the map $\pi^{++}{\restriction}D_{i}$ is a bijection,
  • $\pi^{++}[D_{1}]=\pi^{++}[D_{8}]=J^{+}$,
  • $\pi^{++}[D_{2}]=\pi^{++}[D_{7}]=C_{3}^{+}$,
  • $\pi^{++}[D_{i}]=K^{+}$ for $i=3,4,5,6$.
@@
(5.35) $(\pi^{++}{\restriction}D_{8}{\textrm{ is an isomorphism with }}J^{+})\ \&\ (d\in D_{8}\wedge dRD_{7}\ \to\pi^{++}(d)\in C_{3}^{+}),$
(5.36) $(\pi^{++}{\restriction}D_{1}{\textrm{ is an isomorphism with }}J^{+})\ \&\ (d\in D_{1}\wedge dRD_{2}\ \to\pi^{++}(d)\in C_{3}^{+}),$
Next, we are going to use $M$ and $f$ from $(7)$ to construct $J^{\prime\prime}$. W.l.o.g. we can assume that $f$ factors through $\pi^{+}\circ\pi^{++}$, or simply $f$ maps onto $J^{++}$, so we can pick $J^{\prime\prime}$, $f^{+}\in{\rm Epi}(J^{\prime\prime},M)$, $\pi^{\prime\prime}\in{\rm Epi}(J^{\prime\prime},J^{++})$ with $\pi^{\prime\prime}=f\circ f^{+}$. We can assume that $(\pi^{\prime\prime})^{-1}(a)$ consists of intervals each of length at least $3$ whenever $a\in J^{++}$.
Finally, we claim that letting $\pi^{\prime}=\pi^{+}\circ\pi^{++}\circ\pi^{\prime\prime}$, $h=h^{+}\circ\pi^{++}\circ\pi^{\prime\prime}$, and $K^{\prime\prime}$ be any interval in $J^{\prime\prime}$ with
  • $\pi^{\prime\prime}[K^{\prime\prime}]=D_{4}$, and $K^{\prime\prime}$ is maximal such interval (that is, $a\notin K^{\prime\prime}$, $aRK^{\prime\prime}$ implies $\pi^{\prime\prime}(a)\in D_{3}\cup D_{5}$),
@@
will work (that is, $\boxplus_{6}$ is satisfied). We also remark that passing to a further extension of $J^{\prime\prime}$, if necessary and replacing $J^{\prime\prime}$, $\pi^{\prime\prime}$, without loss of generality such $K^{\prime\prime}$ exists.
@@
It is easy to see that $\pi^{\prime\prime}(c)\in D_{3}\cup D_{5}$. First we assume that $\pi^{\prime\prime}(c)\in D_{3}$, and we let $a$ denote the unique node in $K^{+}$ with $aRC_{2}^{+}$. By the way $\pi^{++}:J^{++}\to J^{+}$ is defined one can check that
$\pi^{++}\circ\pi^{\prime\prime}(c)=a.$
@@
(5.37) ${\rm tp}^{c,C}(\pi^{++}\circ\pi^{\prime\prime})\supseteq\{\{a\},K^{+},K^{+}\cup C_{3}^{+},K^{+}\cup C_{3}^{+}\cup C_{2}^{+},K^{+}\cup C_{3}^{+}\cup C_{2}^{+}\cup L^{+}\}.$
Clearly ${\rm tp}^{c,C}(\pi^{\prime\prime})$ contains a set $H$, such that $H$ is either the entire $D_{3}$ and a subset of $D_{4}\cup D_{5}\cup D_{6}$, or $H$ contains the entire $D_{6}$, and a subset of $D_{3}\cup D_{4}\cup D_{5}$. Either case $\pi^{++}[H]=K^{+}$. Similarly, ${\rm tp}^{c,C}(\pi^{\prime\prime})$ contains a set $H^{\prime}$ that is either the union of $D_{2}\cup D_{3}$ and a subset of $D_{4}\cup D_{5}\cup D_{6}\cup D_{7}$, or $H$ is the union of the full $D_{6}\cup D_{7}$ and a subset $D_{2}\cup D_{3}\cup D_{4}\cup D_{5}$. In either case, $\pi^{++}[H]=K^{+}\cup C_{3}^{+}$. This implies that
${\rm tp}^{c,C}(\pi^{++}\circ\pi^{\prime\prime})\supseteq\{\{a\},K^{+},K^{+}\cup C_{3}^{+}\}.$
Now recalling (5.35) and (5.36) it is easy to check that (5.37) holds, indeed. Finally, $k^{+}_{-}$ is an endpoint of $K^{+}$, so by $\blacksquare_{1}$ and $\blacksquare_{3}$ the node $\pi^{+}(k^{+}_{-})=\pi^{\prime}(c)=h^{+}(k^{+}_{-})=h^{\prime}(c)$ is an endpoint of $K^{\prime}$, so applying $\pi^{+}$ and $h^{+}$ to (5.37),
${\rm tp}^{c,C}(\pi^{+}\circ\pi^{++}\circ\pi^{\prime\prime})\supseteq\{\{\pi^{+}(a)\},K^{\prime},K^{\prime}\cup C_{3}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime}\cup L^{\prime}\},$
@@
${\rm tp}^{c,C}(h^{+}\circ\pi^{++}\circ\pi^{\prime\prime})\supseteq\{\{\pi^{+}(a)=h^{+}(a)\},K^{\prime},K^{\prime}\cup C_{3}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime},K^{\prime}\cup C_{3}^{\prime}\cup C_{2}^{\prime}\cup L^{\prime}\},$
which by our assumption $(4)$ on the structure of $J^{\prime}$ uniquely determine ${\rm tp}^{c,C}(\pi^{+}\circ\pi^{++}\circ\pi^{\prime\prime})$ and ${\rm tp}^{c,C}(h^{+}\circ\pi^{++}\circ\pi^{\prime\prime})$, in particular these two coincide.
On the other hand, if $\pi^{\prime\prime}(c)\in D_{5}$, and we let $b$ be the unique node in $K^{+}$ with $bRC_{3}^{+}$, then the same argument shows
(5.38) ${\rm tp}^{c,C}(\pi^{++}\circ\pi^{\prime\prime})\supseteq\{\{b\},K^{+},K^{+}\cup C_{3}^{+},K^{+}\cup C_{3}^{+}\cup C_{2}^{+},K^{+}\cup C_{3}^{+}\cup C_{2}^{+}\cup L^{+}\},$
to which one can similarly apply $h^{+}$ and $\pi^{+}$, and conclude
${\rm tp}^{c,C}(\pi^{+}\circ\pi^{++}\circ\pi^{\prime\prime})={\rm tp}^{c,C}(h^{+}\circ\pi^{++}\circ\pi^{\prime\prime}).$
Appendix A Projective Fraïssé theory
We recall here the framework of projective Fraïssé theory. The theory was introduced in [9]. The description below is a generalized version of [9]. Propositions A.2A.5 are new.
@@ 2. $({\rm G}2)$
The following proposition is useful in applications.
Proposition A.2.
Conditions (i)–(iii) below are equivalent to each other.
  1. (i)
    For each non-empty clopen set $V\subseteq{\mathbb{K}}$, there is $x\in V$ such that $R^{\mathbb{K}}(x)\subseteq V$.
  2. (ii)
    The set $\{x\in{\mathbb{K}}\mid R^{\mathbb{K}}(x)=\{x\}\}$ is dense in $\mathbb{K}$}.
  3. (iii)
    For each clopen set $V\subseteq{\mathbb{K}}$, the set $R^{\mathbb{K}}(V)\setminus V$ is nowhere dense in $\mathbb{K}$}.
Condition (iv) below implies conditions (i)–(iii).
  1. (iv)
    for all $a\in A\in{\mathcal{K}}$, there exist $b\in B\in{\mathcal{K}}$ and an epimorphism $f:B\to A$ in $\mathcal{K}$} such that $f(R^{B}(b))=\{a\}$.
Proof.
We start by showing that conditions (i)–(iii) are equivalent to each other. These arguments use only the fact that $R^{\mathbb{K}}$ is a compact symmetric relation on $\mathbb{K}$}.
[…]
$\{x\in{\mathbb{K}}\mid{\rm diam}(R^{\mathbb{K}}(x))\geq r\}$
is closed. So, by the Baire category theorem, there exists a non-empty clopen set $U^{\prime}\subseteq U$ and $r>0$ such that ${\rm diam}(R^{\mathbb{K}}(x))\geq r$ for all $x\in U^{\prime}$. Let $V\subseteq U^{\prime}$ be a non-empty clopen set with ${\rm diam}(V)<r$. It is clear that (i) fails for this $V$. The implication is proved.
[…]
$R^{{\mathbb{K}}}({\mathbb{K}}\setminus V)\cap V=R^{{\mathbb{K}}}({\mathbb{K}}\setminus V)\setminus({\mathbb{K}}\setminus V)\not\supseteq V.$
So we have $R^{{\mathbb{K}}}({\mathbb{K}}\setminus V)\not\supseteq V$. Any point $x\in V\setminus R^{{\mathbb{K}}}({\mathbb{K}}\setminus V)$ satisfies (i) since $R^{\mathbb{K}}$ is symmetric.
We now show that (iv) implies conditions (i)–(iii).
[…]
$V\supseteq\{x\in{\mathbb{K}}\mid\phi(x)=a\}.$
Let $b\in B\in{\mathcal{K}}$ and an epimorphism $f\colon B\to A$ in $\mathcal{K}$} be chosen for $a$ and $A$ as in point (iv). Now, let $\psi\colon{\mathbb{K}}\to B$ be an epimorphism in $\mathcal{K}$$^{*}$ such that
$\phi=f\circ\psi.$
Pick $x\in{\mathbb{K}}$ such that $\psi(x)=b$. We claim that $x$ is as required by the conclusion of (i). Let $y\in{\mathbb{K}}$ be such that $xR^{\mathbb{K}}y$. We need to show that $y\in V$, for which it is enough to prove that $\phi(y)=a$. We have
$\psi(x)R^{B}\psi(y)\;\hbox{ and }\;\psi(x)=b,$
so, by the choice of $f$,
$f(\psi(y))=a.$
Since $f(\psi(y))=\phi(y)$, it follows that $\phi(y)=a$, and the conclusion follows. ∎
@@
be the quotient map, which we call the canonical projection.
Proposition A.3.
Assume $\mathcal{K}$} is a countable transitive projective Fraïssé class. If $x\in{\mathbb{K}}$, then sets of the form
@@
Suppose that $U\subseteq K$ is open and ${\rm pr}(x)\in U$. Since there are only countably many epimorphisms in $\mathcal{K}$}, by the amalgamation property of $\mathcal{K}$} (and by $({\rm P}1)$, $({\rm P}2)$), we can enumerate a cofinal system $(\varphi_{i})_{i}$ of epimorphisms, in the sense that each epimorphism factors through all but finitely many $\varphi_{i}$.
Suppose that $U$ does not contain any subset of the prescribed form, in particular for each $\varphi_{i}$ there exists $y_{i}\in{\mathbb{K}}$ with ${\rm pr}(y_{i})\notin U$, $\varphi_{i}(y_{i})R\varphi_{i}(x)$. Since every epimorphism from $\mathbb{K}$} into an object in $\mathcal{K}$} factors through a $\varphi_{j}$, w.l.o.g. for each $i\leq j$ there exists $\pi$ such that $\varphi_{i}=\pi\circ\varphi_{j}$, and by a standard compactness argument we can assume $\varphi_{i}(y_{j})=\varphi_{i}(y_{i}))$ for $i\leq j$. By compactness of $\mathbb{K}$}, we can assume that $(y_{i})_{i}$ converges to some $y\in{\mathbb{K}}$ (in fact the convergence of $(\varphi_{i}(y_{j}))_{j}$ already implies this), so by continuity, for each $i$,
$\varphi_{i}(y)=\varphi_{i}(\lim_{j}(y_{j}))=\lim_{j}(\varphi_{i}(y_{j}))=\varphi_{i}(y_{i}),$
@@
Proposition A.4.
Assume $\mathcal{K}$} is a countable transitive projective Fraïssé class. Assume that $\mathcal{K}$} fulfills condition (iv) (or just conditions (i)–(iii)) of Proposition A.2. Let $\phi\colon{\mathbb{K}}\to A$ be an epimorphism in $\mathcal{K}$} and let $S\subseteq A$. Then
  1. (i)
    ${\rm int}\big({\rm pr}\big(\phi^{-1}(S)\big)\big)=K\setminus{\rm pr}\big(\phi^{-1}(A\setminus S)\big)$;
  2. (ii)
    ${\rm pr}\big(\phi^{-1}(S)\big)$ is a regular closed subset of $K$.
Proof.
Set $D_{S}=\phi^{-1}(S)$.
(i) We have
(A.1) ${\rm int}({\rm pr}(D_{S}))\cap{\rm pr}(D_{A\setminus S})=\emptyset.$
Otherwise there is $z\in D_{A\setminus S}$ with ${\rm pr}(z)\in{\rm int}({\rm pr}(D_{S}))$. By continuity of $\rm pr$, if $U\subseteq{\mathbb{K}}$ is a small enough clopen set containing $z$, then
$\emptyset\not=U\subseteq D_{A\setminus S}\;\hbox{ and }\;{\rm pr}(U)\subseteq{\rm pr}(D_{S}).$
So, for this clopen set $U$, we have
$\emptyset\not=U\subseteq D_{A\setminus S}\;\hbox{ and }\;R^{{\mathbb{K}}}(y)\cap D_{S}\not=\emptyset,\hbox{ for each }y\in U.$
Since $D_{S}\cap D_{A\setminus S}=\emptyset$, this implies
$R^{\mathbb{K}}({\mathbb{K}}\setminus U)\supseteq U\not$$\emptyset$,
which, with $V={\mathbb{K}}\setminus U$, contradicts condition (iii) of Proposition A.2. We get (i), that is,
${\rm int}({\rm pr}(D_{S}))=K\setminus{\rm pr}(D_{A\setminus S}),$
with the inclusion $\subseteq$ being a rephrasing of (A.1) and the inclusion $\supseteq$ following from $D_{S}\cup D_{A\setminus S}={\mathbb{P}}$, continuity of the function $\rm pr$, and compactness of $D_{A\setminus S}$.
(ii) Applying (i) to obtain the first and third equalities, we get
${\rm cl}\big({\rm int}({\rm pr}(D_{S}))\big)={\rm cl}(K\setminus{\rm pr}(D_{A\setminus S}))=K\setminus{\rm int}({\rm pr}(D_{A\setminus S}))={\rm pr}(D_{S}).$
@@
We register the following proposition.
Proposition A.5.
Let $\mathcal{K}$} be a countable transitive projective Fraïssé class. Assume $\mathcal{K}$} fulfills condition (iv) (or just conditions (i)–(iii)) of Proposition A.2. Then the canonical homomorphism ${\rm pr}\colon{\rm Aut}({\mathbb{K}})\to{\rm Homeo}(K)$ is injective.
Proof.
We will use condition (ii) of Proposition A.2. Let $f,g\in{\rm Aut}({\mathbb{K}})$ be such that ${\rm pr}(f)={\rm pr}(g)$. Let
$T=\{x\in{\mathbb{K}}\mid R^{\mathbb{K}}(x)=\{x\}\}.$
Since $f$ and $g$ are automorphisms of $({\mathbb{K}},R^{\mathbb{K}})$, we see that $f(T)=g(T)=T$. So, by the assumption ${\rm pr}(f)={\rm pr}(g)$ and condition (A.2), we have that $f(x)=g(x)$ for all $x\in T$. Since, by our assumption (ii), $T$ is dense in $\mathbb{K}$}, we get $f=g$. ∎